OpenStudy (anonymous):

A colony that initially has 70 bacteria triples every day. Write a natural growth model to express the number of bacteria in the colony as a function of time in days.

OpenStudy (amistre64):

A{0} = 70; A{n} = 3(A{n-1}) is our difference equation

OpenStudy (amistre64):

A{1} = 3(70) A{2} = 3(3(70)) A{3} = 3(3(3(70))) = 3^3 (70) A{4} = 3^4 (70) A{n} = 3^n (70); should do it

OpenStudy (anonymous):

this is what I found on this matter (don't remember any more... ): http://en.wikipedia.org/wiki/Exponential_growth

OpenStudy (anonymous):

Thank You

OpenStudy (amistre64):

yep